OpenStudy (anonymous):

Find two geometric means between 7 and 875.

OpenStudy (amilapsn):

\[\Large\sf{7,*,*,875}\]

OpenStudy (anonymous):

I know that much. But what do I do next?

jimthompson5910 (jim_thompson5910):

In the first blank will go 7r where r is the common ratio in the next blank will go 7r*r = 7r^2 agreed?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

the next term after 7r^2 is 7r^2*r = 7r^3 we know that term is 875, so, 7r^3 = 875

jimthompson5910 (jim_thompson5910):

solve 7r^3 = 875 for r and use that to compute 7r and 7r^2

OpenStudy (anonymous):

is it 35 and 175?

jimthompson5910 (jim_thompson5910):

very good